OpenStudy (anonymous):

A projectile is fired with an initial speed of 200 m/s, angle of elevation 60 degrees, and fired from 100m above the ground. What is the range, maximum height, and speed at impact?

OpenStudy (anonymous):

Fundamental kinematics equation ought to take care of this one: \[\Delta s=vt+0.5at^2\]

OpenStudy (anonymous):

I recommend treating the horizontal and vertical directions separately by taking vector components as well.

OpenStudy (anonymous):

We haven't really talked much about physics in this class (multivariable) and I haven't taken basic physics for awhile. What would be acceleration here?

OpenStudy (cwrw238):

at maximum height velocity v = 0 vertical component of velocity at start is 200*sin 60 m/s use v^2 = u^2 + 2as where s = max height, and a = -9.81m s^-2 so 0 = (200*sin60) ^2 - 2*9.81* s solve this for s

OpenStudy (anonymous):

If we're assuming that the projectile is pretty near to Earth's surface, you can use -9.81 m/s/s as gravitational field strength (acceleration).

OpenStudy (cwrw238):

acceleration = gravity = - 9.81 m s_2

OpenStudy (anonymous):

Also, always draw a picture!